Bus tickets from Toledo to Durham are 26 dollars for adults and 18 for students. How many students were there is 1668 was collected from 70 passengers.
I have the answer key but not sure of the equation..trying to help son

Answer by Cromlix(4381)

x = adults
y = students
x + y = 70 ......1
26x + 18y = 1668.......2
Multiply (1) by 18
18x + 18y = 1260....3
26x + 18y = 1668....2
Subtract (3) from (2)
8x = 408
x = 51
Substitute x = 51 into (1)
x + y = 70.....1
51 + y = 70
y = 70 - 51
y = 19
Therefore there were:
51 adults and 19 students.
Hope this helps.
:-)